### v4

@@ I Introduction
The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[1], largely driven by the developments of deep learning algorithms. Despite these advancements, traditional artificial neural networks (ANNs) often fall short in modeling the intricate and dynamic nature of biological neural systems[2]. To bridge this gap, spiking neural networks (SNNs) have emerged as a promising alternative, leveraging the temporal dimension of neuronal activity to more closely emulate the behavior of biological brains.
SNNs utilize discrete spike events to transmit information between neurons, mirroring the asynchronous and event-driven communication observed in biological neurons[3]. This approach not only enhances the biological plausibility of neural models but also offers potential advantages in terms of computational efficiency and power consumption, particularly for hardware implementations[4]. The inherent sparsity and temporality in coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [5].
Despite their potential, SNNs pose significant challenges, particularly in training methodologies and network optimization. Traditional backpropagation techniques used in ANNs are not directly applicable to SNNs due to their discontinuous and non-differentiable nature[6]. Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [7, 8] and surrogate gradient methods [9, 10, 11], have been developed to address these challenges, enabling the effective training of SNNs.
Prior research in ANN has explored the application of bit planes (e.g. [12]) and color models [13, 14] to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing. Motivated by these advancements, we aim to investigate whether bit planes and color models can similarly improve the performance of SNNs. This led us to formulate our research questions:
@@
A preliminary version of this paper is published in [15], where the benefit of using bit plane coding is proved by three grayscale datasets. In this extended version, the following new points have been added.
@@
$L$ A list-like [16] data structure, used to temporarily store coding result
@@ II-A Overview of SNNs
SNNs can be considered a biologically inspired variant of ANNs that offer a different computational paradigm. Unlike traditional ANNs, which process information in a continuous manner using floating-point activations and backpropagation-based learning [17], SNNs rely on discrete events called spikes to encode and transmit information [18]. This event-driven nature of SNNs allows for energy-efficient computations, making them suitable for neuromorphic hardware implementations [19]. While the majority of mainstream ANNs architecture operate on floating point data representations, SNNs mimic the sparse, time-dependent processing found in biological neurons [20]. Furthermore, SNNs employ temporal dynamics to learn and process temporal information naturally, leveraging mechanisms such as STDP for unsupervised learning [21]. Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [9, 11, 22], hybrid SNN training [23] or ANN-to-SNN conversion approaches [24]. Thus, SNNs present a compelling alternative to ANNs for energy-efficient and biologically plausible computations, while also introducing unique challenges in model design and training.
One of the earliest neuron models that is still widely applied in modern SNN architectures is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [25]. In this model, the neuron is characterized by its membrane potential $V$:
@@
While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons[26, 27, 28], challenges still remain in the conversion process. One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches. Despite the introduction of both data-driven and model-driven threshold determination methods[29], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts. Additionally, IF neurons lack time-dependent memory. Once a neuron receives sub-threshold potentials, the membrane potential persists without decay until the neuron fires. This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [30].
Still, recent study have shown that SNN employed with IF neurons are more energy efficient than traditional ANN[31]. Within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better than leaky integrate-and-fire (LIF) neurons[30], with the SEW-ResNet[11] architecture being one such example.
Figure 1: Visualization of the extracted bit planes from a sample image in the CIFAR-10 dataset [32]. Higher-order bit planes (e.g., most significant bits) encode more spatial and structural information, capturing the prominent features of the image. In contrast, lower-order bit planes (e.g., least significant bits) primarily contain finer details and noise, contributing less to the overall image structure.
@@ II-B Deep Residual Learning in SNNs
The success of ResNet[33] in deep learning has motivated efforts to extend residual learning to SNN. Prior attempts, commonly referred to as Spiking ResNets[28], have largely replicated the architecture of ResNet[33] by replacing ReLU activation functions with spiking neurons. However, this approach presents certain limitations:
  • Inability to achieve identity mapping across neuron models: Residual learning is predicated on the concept of identity mapping, where layers can be trained to approximate an identity function. While this is straightforward in ResNets[33] with ReLU activation, it is not always feasible in Spiking ResNets[28]. Specifically, [11] claimed that for certain spiking neuron models, such as LIF neurons[25] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging. This complicates the realization of identity mapping.
  • Vanishing/exploding gradients: Even in cases where identity mapping can be achieved, Spiking ResNets[28] are prone to vanishing or exploding gradients during training. This occurs because the gradient propagation through multiple layers of spiking neurons may either diminish to zero or grow uncontrollably, complicating the training of deeper networks.
To address these issues, SEW ResNet[11] was introduced as a replacement design for residual blocks while utilizing IF neurons. The SEW block capitalizes on the binary nature of spikes by applying element-wise functions $g$ to the output of the residual mapping $A_{l}[t]$ and the input spike train $S_{l}[t]$[11]. This approach facilitates identity mapping, mitigates the vanishing/exploding gradient problem and achieves higher accuracy with fewer time steps compared to existing SNN models.
In recent researches, using backpropagation methods to train SNNs [34, 10, 9, 11] has become a common practice in neuromorphic computing. In SNNs, spike signals are typically modeled using binary values (0s and 1s) from a Heaviside function $H(\theta,t)$, where $\theta$ is the membrane potential and $t$ is a threshold:
@@
This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively [11].
Similar to other new types of neural networks [35, 36, 37], the choice of input data representation may have strong impacts on output performance. Thus, in this paper we investigate a new way that employs bit plane coding for data representation in SNNs.
@@ III-A Bit planes in deep learning
Bit planes have shown extensive utility in deep learning applications, offering new ways to improve model efficiency and performance. In [12], researchers used bit planes to encode the input layer in Binary Neural Network (BNN). They argued that bit planes retain crucial spatial information, enabling a reduction in the number of multiply-accumulate operations (MACs) without compromising the model’s accuracy or robustness. This demonstrates the potential of bit planes in streamlining computations while maintaining competitive performance.
The use of bit plane slicing on RGB images for adversarial robustness was explored in [38]. By selecting the most significant bit planes, the authors developed a classification model that was better equipped to handle adversarial attacks. This approach capitalizes on the spatial information encoded in different bit planes, which helps the model resist manipulations that would otherwise degrade its performance.
In the domain of medical imaging, bit plane slicing has also been applied to improve accuracy in breast cancer recognition tasks[39]. Researchers found that lower-order bit planes are more susceptible to noise, and by removing them before inputting the data into a convolutional neural network (CNN), they achieved higher classification accuracy. This method highlights how bit plane analysis can be tailored to specific challenges, such as noise reduction in medical images, to enhance model effectiveness.
@@ III-B Color models in deep learning
Color models are fundamental in computer graphics as they define how colors are represented, manipulated, and displayed for digital images. In computer graphics applications, understanding and effectively managing color models is crucial for tasks such as rendering realistic scenes, creating digital artwork, and designing user interfaces. Color models can be divided into device-oriented models, user-oriented models and device-independent models [40, 41].
@@
User-oriented color models are considered as a pathway between the observer and the device handling the color information. For our study, we mostly investigate cylindrically described models, including HSL and HSV. According to [40], these models also inherit other special properties such as device dependence and cylindrical-coordinate representation.
@@
Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks, such as image recognition[42, 13], image segmentation[14] and adversarial attacks prevention[43].
@@ IV Proposed Method
To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[44], which is a simple dataset containing grayscale images of hand-written numbers. We compared three approaches: traditional rate coding, bit plane coding only, and the combination of both rate coding and bit plane coding. The detailed settings of this experiment can be found in Section VI-A. The results are provided in Table III.
Table III: Average validation accuracy comparison on MNIST[44]
@@
As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding. Initially, the process commences with the conventional spike coding procedure applied to the image tensor. The bit plane coding algorithm is then employed to transform an input image into binary planes. The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from spike coding and bit plane coding. These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes.
@@
function BitplaneEncode($X,X_{max}$, colormodelcolormodel)
@@ V Theoretical basis
As discussed in Section II-B, surrogate gradient methods attempt to adapt deep learning’s gradient optimization techniques to SNNs, maintaining a close connection to traditional deep learning and convex optimization. A classical stochastic optimization problem associated with learning a target function $h$ can be formulated as[45]:
@@
where $\ell$ is a loss function, $\mathbf{x}$ represents stochastic inputs (assumed to be vectors in Euclidean space), and $p_{\mathbf{w}}$ is a differentiable predictor parameterized by $\mathbf{w}$. It is assumed that $E_{x}[\left\|\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}\right\|^{2}]\leq G(\mathbf{w})^{2}$ for some scalar $G(\mathbf{w})$, and that $F$ is differentiable[45, 46].
@@
Previous studies in gradient-based optimization[45, 46] have demonstrated a positive correlation between a model’s performance and the signal-to-noise ratio (SNR) of the stochastic gradient, defined by:
@@ VI-A Experimental settings
To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [11]. The model and experiment process were implemented using PyTorch [47] and SpikingJelly[48]. Color conversion algorithms are based on implementations from Kornia [49]. For the baseline method in the experiments, the Poisson-distributed rate coding [50, 51] is employed with a step size $T=10$ where spike $s_{T,i}$ for i-th pixel at step $T$ with intensity $P_{T,i}$ can be interpreted as:
$s_{T,i}=\begin{cases}1&\text{if }random(0.0,1.0)\leq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (12)
To enable efficient training, surrogate gradient methods were applied to all IF neurons. We utilized the arctangent surrogate gradient function S(U)S(U)[11, 10]. Also, we adopted the standard cross-entropy loss function with a softmax activation, commonly adopted for classification tasks[47, 52, 53], computed using mean reduction across mini-batches.
Inspired by traditional deep learning approach in model fine-tunning [52, 53], the optimization process is conducted with the Adam algorithm [54], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets.
In the first part of our experiments, we only use grayscale image datasets to clearly see the effects of bit plane coding. In the second part, popular datasets of color images are used to see the practical effects of both bit plane coding and color models. We allocated 80% of each dataset for training and the remaining 20% with deterministic random seeding for validation. All datasets were processed with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU.

### v5

@@ I Introduction
The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[wani2020advances], largely driven by the developments of deep learning algorithms. Despite these advancements, traditional artificial neural networks (ANNs) often fall short in modeling the intricate and dynamic nature of biological neural systems[nguyen2021review]. To bridge this gap, spiking neural networks (SNNs) have emerged as a promising alternative, leveraging the temporal dimension of neuronal activity to more closely emulate the behavior of biological brains.
SNNs utilize discrete spike events to transmit information between neurons, mirroring the asynchronous and event-driven communication observed in biological neurons[auge2021survey]. This approach not only enhances the biological plausibility of neural models but also offers potential advantages in terms of computational efficiency and power consumption, particularly for hardware implementations[blouw2019benchmarking]. The inherent sparsity and temporality in coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [rajendran2019low].
Despite their potential, SNNs pose significant challenges, particularly in training methodologies and network optimization. Traditional backpropagation techniques used in ANNs are not directly applicable to SNNs due to their discontinuous and non-differentiable nature[nunes2022spiking]. Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [srinivasan2017spike, liu2021sstdp] and surrogate gradient methods [neftci2019surrogate, eshraghian2023training, fang2021deep], have been developed to address these challenges, enabling the effective training of SNNs.
Prior research in ANN has explored the application of bit planes (e.g. [vorabbi2023input]) and color models [gowda2019colornet, taipalmaa2020different] to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing. Motivated by these advancements, we aim to investigate whether bit planes and color models can similarly improve the performance of SNNs. This led us to formulate our research questions:
@@
A preliminary version of this paper is published in [luu2024], where the benefit of using bit plane coding is proved by three grayscale datasets. In this extended version, the following new points have been added.
@@
$L$ A list-like [abelson1996structure] data structure, used to temporarily store coding result
@@ II-A Overview of SNNs
SNNs can be considered a biologically inspired variant of ANNs that offer a different computational paradigm. Unlike traditional ANNs, which process information in a continuous manner using floating-point activations and backpropagation-based learning [rumelhart1986learning], SNNs rely on discrete events called spikes to encode and transmit information [maass1997networks]. This event-driven nature of SNNs allows for energy-efficient computations, making them suitable for neuromorphic hardware implementations [merolla2014million]. While the majority of mainstream ANNs architecture operate on floating point data representations, SNNs mimic the sparse, time-dependent processing found in biological neurons [gerstner2002spiking]. Furthermore, SNNs employ temporal dynamics to learn and process temporal information naturally, leveraging mechanisms such as STDP for unsupervised learning [lee2018training]. Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [neftci2019surrogate, fang2021deep, luu2025accuracy], hybrid SNN training [luu2025hybrid] or ANN-to-SNN conversion approaches [rueckauer2017conversion]. Thus, SNNs present a compelling alternative to ANNs for energy-efficient and biologically plausible computations, while also introducing unique challenges in model design and training.
One of the earliest neuron models that is still widely applied in modern SNN architectures is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [abbott1999lapicque]. In this model, the neuron is characterized by its membrane potential $V$:
@@
While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons[li2021free, deng2021optimal, hu2021spiking], challenges still remain in the conversion process. One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches. Despite the introduction of both data-driven and model-driven threshold determination methods[diehl2015fast], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts. Additionally, IF neurons lack time-dependent memory. Once a neuron receives sub-threshold potentials, the membrane potential persists without decay until the neuron fires. This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [zhang2019tdsnn].
Still, recent study have shown that SNN employed with IF neurons are more energy efficient than traditional ANN[dampfhoffer2022snns]. Within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better than leaky integrate-and-fire (LIF) neurons[zhang2019tdsnn], with the SEW-ResNet[fang2021deep] architecture being one such example.
Figure 1: Visualization of the extracted bit planes from a sample image in the CIFAR-10 dataset [alex2009learning]. Higher-order bit planes (e.g., most significant bits) encode more spatial and structural information, capturing the prominent features of the image. In contrast, lower-order bit planes (e.g., least significant bits) primarily contain finer details and noise, contributing less to the overall image structure.
@@ II-B Deep Residual Learning in SNNs
The success of ResNet[he2016deep] in deep learning has motivated efforts to extend residual learning to SNN. Prior attempts, commonly referred to as Spiking ResNets[hu2021spiking], have largely replicated the architecture of ResNet[he2016deep] by replacing ReLU activation functions with spiking neurons. However, this approach presents certain limitations:
  • Inability to achieve identity mapping across neuron models: Residual learning is predicated on the concept of identity mapping, where layers can be trained to approximate an identity function. While this is straightforward in ResNets[he2016deep] with ReLU activation, it is not always feasible in Spiking ResNets[hu2021spiking]. Specifically, [fang2021deep] claimed that for certain spiking neuron models, such as LIF neurons[abbott1999lapicque] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging. This complicates the realization of identity mapping.
  • Vanishing/exploding gradients: Even in cases where identity mapping can be achieved, Spiking ResNets[hu2021spiking] are prone to vanishing or exploding gradients during training. This occurs because the gradient propagation through multiple layers of spiking neurons may either diminish to zero or grow uncontrollably, complicating the training of deeper networks.
To address these issues, SEW ResNet[fang2021deep] was introduced as a replacement design for residual blocks while utilizing IF neurons. The SEW block capitalizes on the binary nature of spikes by applying element-wise functions $g$ to the output of the residual mapping $A_{l}[t]$ and the input spike train $S_{l}[t]$[fang2021deep]. This approach facilitates identity mapping, mitigates the vanishing/exploding gradient problem and achieves higher accuracy with fewer time steps compared to existing SNN models.
In recent researches, using backpropagation methods to train SNNs [huh2018gradient, eshraghian2023training, neftci2019surrogate, fang2021deep] has become a common practice in neuromorphic computing. In SNNs, spike signals are typically modeled using binary values (0s and 1s) from a Heaviside function $H(\theta,t)$, where $\theta$ is the membrane potential and $t$ is a threshold:
@@
This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively [fang2021deep].
Similar to other new types of neural networks [ngo2023survey, nguyen2022evaluation, nguyen2022quantum], the choice of input data representation may have strong impacts on output performance. Thus, in this paper we investigate a new way that employs bit plane coding for data representation in SNNs.
@@ III-A Bit planes in deep learning
Bit planes have shown extensive utility in deep learning applications, offering new ways to improve model efficiency and performance. In [vorabbi2023input], researchers used bit planes to encode the input layer in Binary Neural Network (BNN). They argued that bit planes retain crucial spatial information, enabling a reduction in the number of multiply-accumulate operations (MACs) without compromising the model’s accuracy or robustness. This demonstrates the potential of bit planes in streamlining computations while maintaining competitive performance.
The use of bit plane slicing on RGB images for adversarial robustness was explored in [liu2022defending]. By selecting the most significant bit planes, the authors developed a classification model that was better equipped to handle adversarial attacks. This approach capitalizes on the spatial information encoded in different bit planes, which helps the model resist manipulations that would otherwise degrade its performance.
In the domain of medical imaging, bit plane slicing has also been applied to improve accuracy in breast cancer recognition tasks[chen2019breast]. Researchers found that lower-order bit planes are more susceptible to noise, and by removing them before inputting the data into a convolutional neural network (CNN), they achieved higher classification accuracy. This method highlights how bit plane analysis can be tailored to specific challenges, such as noise reduction in medical images, to enhance model effectiveness.
@@ III-B Color models in deep learning
Color models are fundamental in computer graphics as they define how colors are represented, manipulated, and displayed for digital images. In computer graphics applications, understanding and effectively managing color models is crucial for tasks such as rendering realistic scenes, creating digital artwork, and designing user interfaces. Color models can be divided into device-oriented models, user-oriented models and device-independent models [ibraheem2012understanding, plataniotis2001color].
@@
User-oriented color models are considered as a pathway between the observer and the device handling the color information. For our study, we mostly investigate cylindrically described models, including HSL and HSV. According to [ibraheem2012understanding], these models also inherit other special properties such as device dependence and cylindrical-coordinate representation.
@@
Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks, such as image recognition[kim2018efficient, gowda2019colornet], image segmentation[taipalmaa2020different] and adversarial attacks prevention[chyad2025exploring].
@@ IV Proposed Method
To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[deng2012mnist], which is a simple dataset containing grayscale images of hand-written numbers. We compared three approaches: traditional rate coding, bit plane coding only, and the combination of both rate coding and bit plane coding. The detailed settings of this experiment can be found in Section VI-A. The results are provided in Table III.
Table III: Average validation accuracy comparison on MNIST[deng2012mnist]
@@
As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding. Initially, the process commences with the conventional spike coding procedure applied to the color converted image tensor (defaulted to RGB if unspecified in our algorithm). The bit plane coding algorithm is then employed to transform an input image into binary planes. The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from spike coding and bit plane coding. These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes.
@@
function BitplaneEncode($X,X_{max}$, $colormodel=\text{"RGB"}$)
@@ V Theoretical basis
As discussed in Section II-B, surrogate gradient methods attempt to adapt deep learning’s gradient optimization techniques to SNNs, maintaining a close connection to traditional deep learning and convex optimization. A classical stochastic optimization problem associated with learning a target function $h$ can be formulated as[shalev2017failures]:
@@
where $\ell$ is a loss function, $\mathbf{x}$ represents stochastic inputs (assumed to be vectors in Euclidean space), and $p_{\mathbf{w}}$ is a differentiable predictor parameterized by $\mathbf{w}$. It is assumed that $E_{x}[\left\|\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}\right\|^{2}]\leq G(\mathbf{w})^{2}$ for some scalar $G(\mathbf{w})$, and that $F$ is differentiable[shalev2017failures, ghadimi2013stochastic].
@@
Previous studies in gradient-based optimization[shalev2017failures, ghadimi2013stochastic] have demonstrated a positive correlation between a model’s performance and the signal-to-noise ratio (SNR) of the stochastic gradient, defined by:
@@ VI-A Experimental settings
To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [fang2021deep]. The model and experiment process were implemented using PyTorch [paszke2019pytorch] and SpikingJelly[doi:10.1126/sciadv.adi1480]. Color conversion algorithms are based on implementations from Kornia [eriba2019kornia]. For the baseline method in the experiments, the Poisson-distributed rate coding [kim2018deep, heeger2000poisson] is employed with a step size $T=10$ where spike $s_{T,i}$ for i-th pixel at step $T$ with intensity $P_{T,i}$ can be interpreted as:
$s_{T,i}=\begin{cases}1&\text{if }\mathcal{N}(0.0,1.0)\leq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (12)
To enable efficient training, surrogate gradient methods were applied to all IF neurons. We utilized the arctangent surrogate gradient function S(U)S(U)[fang2021deep, eshraghian2023training]. Also, we adopted the standard cross-entropy loss function with a softmax activation, commonly adopted for classification tasks[paszke2019pytorch, radford2021learning, luu2023blind], computed using mean reduction across mini-batches.
Inspired by traditional deep learning approach in model fine-tunning [radford2021learning, luu2023blind], the optimization process is conducted with the Adam algorithm [kingma2014adam], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets.
In the first part of our experiments, we only use grayscale image datasets to clearly see the effects of bit plane coding. In the second part, popular datasets of color images are used to see the practical effects of both bit plane coding and color models. We allocated 80% of each dataset for training and the remaining 20% with deterministic random seeding [^1] Due to resource constraints, common validate tactics (such as utilizing k-fold for cross validation, etc.) might not be implementable on deep SNN training, where simulation of multiple time step is required and training cost are commonly higher than ANN on GPU devices. Enforcing deterministic random seed offer a good trade off for accuracy validation with certain level of bias in training and validating process. for validation. All datasets were processed with a batch size of 16 images [^2] While our main purpose for using small batch size is due to resource constraints, prior work have shown that model training with small batch size can still yield optimal solution given proper setting implemented[masters2018revisiting]. for 100 epochs on an NVIDIA RTX 3090 GPU.